# Survey of Load-Altering Attacks Against Power Grids: Attack Impact, Detection and Mitigation

### v1

@@ II-A Attacker Model
In conclusion, the threat model for cyberattacks on power systems revolves around the exploitation of vulnerabilities across both grid-end and consumer-end components. While traditional attacks on centralized systems such as SCADA and AGC are well-established, the rise of DERs and smart IoT devices introduces new entry points for adversaries. These vulnerabilities, exacerbated by complex supply chains and limited security in consumer devices, allow attackers to launch sophisticated, multi-layered LAAs. The attacker can leverage partial system knowledge to modify control actions of IoT and DER devices, potentially destabilizing the grid. Realistically, the attacker’s knowledge is constrained, making statistical, adversarial, and RL models relevant for capturing the nuanced threat dynamics. These models highlight that, despite the limitations in system knowledge, attackers can still inflict significant harm by exploiting communication and control interfaces in IoT-connected devices, thereby requiring robust cybersecurity measures at both the consumer and grid levels.
@@ III-A Impact on Power Grid Operation and Stability
Understanding the impact of LAAs on grid operations and identifying the most disruptive attack points are crucial for risk analysis. Table II summarizes the existing studies on this topic. This table highlights the models studied, and the attack impacts considered in each work.
@@ III-A1 Mechanism of Grid Operations
An unforeseen load surge or drop can cause frequency deviations. The primary frequency response, provided by the system’s inertia and fast-acting technologies like battery storage, compensates for these deviations in the first few seconds of an incident. Secondary and tertiary frequency responses are additional components of frequency control schemes that operate more slowly and are supported by resources like automatic generation control (AGC), spinning reserves, mobile batteries, and off-grid DERs. In distribution systems, such load changes can result in voltage constraint violations. To mitigate this, devices such as shunt capacitors, batteries, flexible AC transmission system (FACTS) devices like soft open points (SOPs), circuit breakers, and DERs can be deployed. During LAAs, the adversary forces sudden, unexpected load changes, potentially disrupting one or both of the discussed equilibria.
@@ III-A2 LAA Impact
To the best of our knowledge, the concept of LAAs was first introduced in [49]. The impact of sudden and large-scale load manipulations, referred to as static LAAs, on transmission systems was first studied using simulation-based approaches. [50] and [51] demonstrated that static LAAs can disrupt the equilibrium between generation and demand, resulting in frequency safety violations. A sudden increase in demand forces the grid operator to rely on primary generator response, potentially leading to increased line power flows. Since the operator lacks control over power flow during the primary response stage, excessive power flows could cause line outages or even cascading failures. Additionally, LAAs may impact steady-state metrics such as operational costs, as operators are compelled to activate more expensive ancillary services to manage increased demand.
@@
Recent studies have also explored how system conditions, particularly low load and low inertia, influence the impact of LAAs. For instance, [53] compared load consumption patterns across seven U.S. cities during the COVID-19 pandemic and demonstrated that LAAs targeting regions with high load proportions have the most severe impact on frequency stability. Motivated by these findings, [24] explored the heightened vulnerability of power grids under low inertia conditions, such as those seen during the pandemic, where errors in RES forecasting can exacerbate the effects of LAAs. Their results showed that under these conditions, attackers can destabilize the system more easily, as low inertia leads to larger frequency swings, compounding the challenges faced by grid operators.
@@
While static LAAs are concerning, dynamic LAAs (DLAAs) represent a more sophisticated and potentially more impactful form of attack. In a DLAA, an attacker repeatedly alters loads in an oscillatory manner, guided by feedback signals from the grid. One variant, as proposed by [54], involves load changes proportional to the local frequency deviation from its nominal value, thus exacerbating the load-generation imbalance in a strategic manner. However, continuous load changes can be difficult to implement in practice. Alternately, [55] and [56] explored DLAAs where load devices are switched on and off in synchronization with inter-area oscillatory frequencies, effectively exciting oscillations and driving the system toward instability.
@@
To further analyze both static and dynamic LAAs, [57] developed an analytical framework leveraging second-order dynamical system theory. Their study derived eigenvalue sensitivities of the grid’s frequency control loop and proposed a method to identify the nodes from which the attacker could launch the most efficient DLAAs. However, many of these works focus on specific LAA scenarios, such as LAAs of defined magnitudes at specific nodes. To address the broader landscape of potential attacks, [58], [59], and [60] employed rare-event sampling techniques to evaluate the impact of LAAs across various combinations of nodes and load magnitudes. Their findings emphasize that DLAAs occurring at discrete intervals, in combination with existing protection measures, can still lead to cascading failures, making dynamic LAAs more disruptive than static ones.
Another emerging area of focus is the impact of LAAs targeting EVCS. In [34], the authors demonstrated that ECVS-based destabilizing LAAs can be launched by solely relying on publicly available information, such as the EVCS location, demand and power grid datasets. They propose a data-driven load manipulation strategy that relocates the system’s eigenvalues to a region of vulnerability (defined by the North American Electric Reliability Corporation standards). Unlike [54, 61, 62, 63], the proposed approach does not require real-time monitoring of the power grid state. In [64], the authors perform a case study of EVCS-based LAAs considering the non-linear nature of the load. They show that the high reactive power demand of EVCS can result in a larger impact on the power grid as compared to LAAs that target residential loads. [65] analyzes the problem of launching DLAAs by coordinating distributed EVCS. Notably, their analysis considers communication delays incurred in sending the load-altering commands as well as outmanoeuvring existing defense solutions that are installed to suppress oscillations in a power system such as the travelling wave or transient energy-based methods. A robust optimization method is implemented to maximize the attack impact by coordinating EVCS from multiple nodes. [66] explores the possibility of exploiting security vulnerabilities of mobile applications used for controlling EV charging infrastructures. They perform real-time simulations on a test bed to show lack of user/vehicle authentication provides security breach opportunities for remote adversaries to gain control over the charging operations and launch attacks by manipulating them.
While most research has focused on LAAs targeting the transmission network, the impact on distribution networks has garnered recent attention. For example, [67] points to issues such as low voltage problems and line overloads in distribution networks due to LAAs and proposes the use of soft open points as the mitigation method (which will be elaborated in Section V). An analytical framework to quantify the impact of LAAs on distribution systems was proposed in [68], which derived closed-form expressions to characterize the minimum required IoT devices to cause voltage safety violations. The work investigated the effect of the location of LAA and voltage-dependency of loads (using the ZIP load model) on the attack impact. It showed that LAAs launched at the downstream nodes of a distribution network (e.g., the leaf nodes) are more impactful. Furthermore, [69] examined the impacts of EVCS-based LAAs EVCS on the voltage profile of distribution networks. Finally, the joint impact of LAAs on both transmission and distribution (T&D) networks has been explored in works such as [70]. Using a bi-level optimization model, the study assessed the socio-economic impacts of LAAs, highlighting metrics like energy not served and increased energy costs. Their findings indicate that the attack’s impact is related to the penetration of IoT devices in the system and the attacker’s strategy, underscoring the importance of understanding the interplay between transmission and distribution networks in mitigating LAA risks.
@@
TABLE II: Overview of the existing works on the impact of LAA on power grid operation and stability.
@@ III-B Impact on Energy Markets
While LAAs pose a serious threat to the stability and operation of power grids, their impact extends beyond grid performance. An increasingly concerning objective of such attacks is economic manipulation [72, 73, 74]. The energy market, which serves as a hub for the substantial financial transactions in the power system, is naturally a focal point for these attackers. By intentionally manipulating demand and supply through compromised devices, attackers can influence market prices, leading to financial gain.
@@ III-B1 Mechanism of Energy Market
The energy market functions as a competitive platform where electricity generators offer their production to meet demand, while electric utilities and other market participants purchase electricity through real-time and day-ahead markets. Fig. 3 illustrates the timeline of electricity market operations, highlighting three key stages: the day-ahead market, the intra-day market, and the real-time market. In the day-ahead market, participants submit bids during the day-ahead bid period, followed by market clearing, where bids are processed, and energy schedules are set [75]. In the intra-day market, participants refine their bids and adjust their positions in response to updated information. The real-time market focuses on real-time flexibility bids, allowing providers to adjust supply and demand [76]. It balances the gaps between day-ahead commitments and the actual electricity demand and supply in real time, which ensures grid stability by aligning supply with real-time demand.
Figure 3: Timeline diagram of the energy markets operation.
Electricity pricing is a fundamental aspect of energy markets. Common pricing schemes include flat pricing, time-of-use (ToU) pricing, and real-time pricing (RTP). While flat pricing sets a constant price throughout the day, ToU pricing categorizes the day into different time periods, each with a specific price that is predetermined and announced before the operation begins [77]. In contrast, RTP adjusts electricity costs based on actual market conditions in real-time, typically on an hourly or sub-hourly basis, reflecting fluctuations in supply and demand. This approach factors in grid conditions, resource availability, and consumption changes. By exposing consumers to real-time price variations, RTP promotes flexible consumption behaviors, enhancing load management and grid stability. It is especially beneficial with the integration of RES, as it aligns consumption with periods of high generation or lower demand, ensuring efficient grid use [78].
@@ III-B2 LAAs to Disrupt the Energy Market Operation
Real-time electricity prices in energy markets are closely tied to system load demands [79], making them susceptible to manipulation through LAAs. By compromising large-scale loads, attackers can influence prices or disrupt the market [80, 81, 82, 83, 84]. Understanding the impact of LAAs on the energy market is therefore crucial, and several studies have shown the feasibility of such attacks. Table III summarizes existing research, which is further discussed below.
TABLE III: Literature of LAAs effect in energy market operations.
@@
A novel IoT attack named Manipulation of Market via IoT (MaMIoT) was presented in [83]. The attack leverages LAAs on high-wattage IoT botnets in order to affect the electricity prices in the energy market. An optimization model is proposed to solve for a stealthy and repeatable LAA using IoT devices, which maximizes the attack’s impact. By increasing the power demand of the compromised devices, it can intentionally increase the electricity price in the real-time market, which in turn enables the power generation utilities to gain more benefits. Results demonstrate that even with a relatively small number of controllable bots, the attacker can significantly increase their profits while inflicting substantial financial losses on other participants in the energy markets.
The authors in [80] investigated the impact on the locational marginal prices (LMPs) in distribution systems when different levels of LAAs are launched in the T&D system. They analyzed how the impacts of LAAs in the targeted distribution systems could propagate through the transmission system to neighboring distribution systems. The results show that by performing LAAs at selected loads, it is possible to make the price of electricity 3×3$\times$3 × higher in specific distribution systems. It also shows that the owners of energy storage systems can utilize LAAs to intentionally create high-loading conditions to sell their energies at higher prices and generate profits.
An LAA against power distribution systems was introduced in [81], which exploited the mutual dependency of the electricity price and the power consumption of the flexible loads in demand response programs. The work analyzed the potential risks of manipulating electricity prices through LAA under a demand response framework. The transactive energy framework was presented to safeguard the power distribution systems against such LAA attacks. By increasing the number of energy transaction cycles per second, the impact of LAAs on system voltage frequency can be mitigated. Moreover, the authors also pointed out that due to malicious changes in electricity prices, electricity consumers may be induced to shift their power loads together to a certain time period, which in turn can destabilize the operation of the power systems. Targeting the demand response in the smart grid as well, authors in [82] studied the potential impacts caused by two kinds of attackers: a malicious attacker who wants to damage the equipment in the power grid and a selfish attacker that wants to benefit from the attack. According to the experimental results, it is possible to largely increase the attackers’ profits by exploiting the direct load control system.
The authors in [84] investigated the charging manipulation attacks (CMAs) against EV charging, which aims to shift the EV aggregator’s demand to different periods of the day. The study delved into the impact of CMAs on aggregators’ profit and performance in the day-ahead markets and real-time markets by modeling their participation in both markets. Results show that the CMAs can lead to a considerable increase in the aggregated surcharge charging cost.
@@ IV Attack Detection and Localization
TABLE IV: Detection and localization approaches against LAAs.
@@
If a successful LAA occurs, swift detection and localization are of utmost importance to limit the damage. Table IV summarizes the work in this area, which can be classified into model-based, data-driven and hybrid approaches. Model-based approaches utilize predefined algorithms or mathematical models to detect and locate LAAs, relying on a thorough understanding of system dynamics. Data-driven approaches utilize ML algorithms to automatically capture potential relationships in the data and anomalies caused by LAAs. Hybrid approaches use a combination of the two.
@@ IV-A Model-Based Approaches
One approach to detect the LAAs is to model the power network as a semi-explicit class of differential-algebraic equations (DAEs) [85] and then use a multi-variable super-twisting sliding mode observer (SMO) to estimate the state of the system. The observer is able to estimate any signals representing failures to detect LAAs, with only one subset of the differential states being measured, i.e., the generator phase angles. Focusing on DLAAs, [86] utilized the estimation of residual signals to detect attacks. Specifically, the power system was first modeled as state-space equations, and then a robust SMO was designed with an observer residual evaluation function. By comparing the value of the function with the ones without attacks, the algorithm identifies whether or not a DLAA is active in the system.
Based on similar ideas of utilizing the observers, the Super-Twisting Sliding Mode Algorithm (STA) was used in the design of LAA detection while also incorporating the usage of BESS in the power system to defend against LAAs [87]. Other work made use of the linear matrix inequality (LMI) to design a bank of residual generators for functional observers to detect LAAs [88]. The approach reduces the number of observed states, which is suitable for calculations in large-scale smart grids. The researchers in [39] designed a closed-loop SMO for the detection of closed-loop DLAAs, whose robustness is enhanced by using the $H_{\infty}$ performance index. A method using a low-rank Kalman filter was proposed in [89], which identifies destabilizing attacks by estimating system states and attack parameters with the rank-1 method, achieving high accuracy and computational efficiency. Authors in [90] proposed using an adaptive fading Kalman filter (AFKF) to estimate the state of the power grid, with which the Gaussian noise can be filtered, and a more accurate state change curve can be obtained. The detection of the DLAAs is based on a Euclidean distance ratio detection algorithm.
Model-based approaches focus on designing observers or state estimators using a state-space model of the system. While these methods provide strong theoretical foundations and require less data when the system is accurately modeled, the increasing integration of diverse equipment into power grids makes accurate modeling more challenging. This issue is particularly evident in large, complex grids, where developing precise state-space models requires significant effort, underscoring a key limitation of model-based approaches.

### v2

@@ II-B Attack Types
In conclusion, vulnerabilities of IoT-enabled load devices are exacerbated due to complex supply chains and limited security in consumer devices, allowing attackers to launch sophisticated and multi-layered LAAs. The attacker can leverage partial system knowledge to modify control actions of IoT and DER devices, potentially destabilizing the grid. Realistically, the attacker’s knowledge is constrained, making statistical, adversarial, and RL models relevant for capturing the nuanced threat dynamics. These models highlight that, despite the limitations in system knowledge, attackers can still inflict significant harm by exploiting communication and control interfaces in IoT-connected devices, thereby requiring robust cybersecurity measures at both the consumer and grid levels.
@@ III-A Impact on Power Grid Operation and Stability
Understanding the impact of LAAs on grid operations and identifying the most disruptive attack points are crucial for risk analysis. Table III summarizes the existing studies on this topic. This table highlights the models studied, and the attack impacts considered in each work.
@@ III-A1 Mechanism of Grid Operations
An unforeseen load surge or drop can cause frequency deviations. The primary frequency response, provided by the system’s inertia and fast-acting technologies like battery storage, compensates for these deviations in the first few seconds of an incident. Secondary and tertiary frequency responses are additional components of frequency control schemes that operate more slowly and are supported by resources like automatic generation control (AGC), spinning reserves, mobile batteries, and off-grid DERs. In distribution systems, such load changes can result in voltage constraint violations. To mitigate this, devices such as shunt capacitors, batteries, flexible AC transmission system (FACTS) devices like soft open points (SOPs), circuit breakers, and DERs can be deployed. During LAAs, the adversary forces sudden, unexpected load changes, potentially disrupting either or both of the discussed equilibria.
@@ III-A2 LAA Impact
To the best of our knowledge, the concept of LAAs was first introduced in [3]. The impact of sudden and large-scale load manipulations, referred to as static LAAs, on transmission systems was first studied using simulation-based approaches. Studies [65] and [68] demonstrated that static LAAs can disrupt the equilibrium between generation and demand, resulting in frequency safety violations. A sudden increase in demand forces the grid operator to rely on primary generator response, potentially leading to increased line power flows. Since the operator lacks control over power flow during the primary response stage, excessive power flows could cause line outages or even cascading failures. Additionally, LAAs can affect steady-state metrics such as operational costs, as operators must activate more expensive ancillary services to manage increased demand.
Not only can attacks on transmission grids cause disruption in them, but also the damage of a well-crafted LAA on the distribution system can be propagated into the transmission grid. According to [78], a naive LAA on the distribution level does not consider the settings of the circuit breaker connecting the distribution system to the transmission grid. As a result, these attacks may trigger the circuit breaker and stop the propagation of the attack impact to the transmission level. However, in an LAA with an insidious strategy, the attacker accounts for the circuit breakers’ settings and confines the resulting additional active and reactive power to stop activating the circuit breaker. This results in the propagation of the attack impact in the connected upstream transmission grid.
@@
Recent studies have also explored how system conditions, particularly low-load and low-inertia circumstances, influence the impact of LAAs. For instance, [66] analyzed load consumption patterns across seven U.S. cities during the COVID-19 pandemic and demonstrated that LAAs targeting regions with high load proportions have the most severe impact on frequency stability. Motivated by these findings, [27] explored the heightened vulnerability of power grids under low inertia conditions, such as those seen during the pandemic, where errors in RES forecasting can exacerbate the effects of LAAs. Their results showed that under these conditions, attackers can destabilize the system more easily, as low inertia leads to larger frequency swings which further complicates grid management for operators.
@@
While static LAAs are concerning, dynamic LAAs (DLAAs) represent a more sophisticated and potentially more impactful form of attack. In a DLAA, an attacker repeatedly alters loads in an oscillatory manner, guided by the feedback signals from the grid. One variant, as proposed by [60], involves load changes proportional to the local frequency deviation from its nominal value, thus exacerbating the load-generation imbalance strategically. However, continuous load changes can be difficult to implement in practice. Alternately, [59] and [77] explored DLAAs where load devices are switched on and off in synchronization with inter-area oscillatory frequencies, effectively exciting oscillations and driving the system toward instability.
@@
To further analyze both static and dynamic LAAs, [64] developed an analytical framework leveraging second-order dynamical system theory. Their study derived eigenvalue sensitivities of the grid’s frequency control loop and proposed a method to identify the nodes from which the attacker could launch the most efficient DLAAs. However, many of these works focus on specific LAA scenarios, such as LAAs of defined magnitudes at specific nodes. To address the broader landscape of potential attacks, [63], [67], and [69] employed rare-event sampling techniques to evaluate the impact of LAAs across various combinations of nodes and load magnitudes. Their findings emphasize that DLAAs occurring at discrete intervals, in combination with existing protection measures, can still lead to cascading failures, rendering dynamic LAAs more disruptive than static ones.
Another emerging focus is on LAAs targeting EVCS. The work in [36] demonstrated that destabilizing EVCS-based LAAs can be executed using publicly available data, such as EVCS locations and grid demand datasets. It proposed a data-driven load manipulation strategy that shifts system eigenvalues to a vulnerable region defined by NERC standards, without requiring real-time grid monitoring, unlike previous works [60, 79, 80, 81]. Sayed et al. [82] conducted a case study highlighting the significant impact of EVCS-based LAAs due to high reactive power demand compared to residential loads. [71] examined coordinated DLAAs using distributed EVCS, factoring in communication delays and countering transient energy-based methods. Their robust optimization strategy maximized attack impact by coordinating multiple EVCS nodes. [83] explored vulnerabilities in mobile apps controlling EV charging, showing through real-time testbed simulations that inadequate vehicle authentication allows remote attackers to manipulate charging operations and launch attacks.
While most papers has focused on LAAs targeting transmission networks, recent work has begun to address their impact on distribution networks. The authors in [76] identified issues like low voltage and line overloads in distribution networks due to LAAs and suggested soft open points as a mitigation strategy (discussed in Section V). Maleki et al. [62] introduced an analytical framework with closed-form expressions to quantify LAA impacts, showing that attacks at downstream nodes (e.g., leaf nodes) are most effective, and highlighted the role of voltage-dependent loads (using the ZIP model). [72] investigated the voltage profile effects of EVCS-based LAAs on distribution networks. Additionally, [78] studied the joint impact on T&D networks using a bi-level optimization model, linking attack impact to IoT penetration and strategies, and noted key metrics like energy not served and increased costs, emphasizing the need to understand T&D interplay for effective LAA risk mitigation.
@@
TABLE III: Overview of the existing works on the impact of LAA on power grid operation and stability.
@@ III-B Impact on Energy Markets
Grid stability can be severely impacted when attackers compromise a large number of loads, resulting in significant frequency deviations and voltage fluctuations that may lead to cascading failures or blackouts. However, the risks posed by LAAs extend beyond grid performance. In addition to disrupting grid operations, LAAs can also be leveraged for economic manipulation, which has emerged as a growing concern [84, 85, 86]. The energy market, being a central component of financial transactions within the power system, is inherently vulnerable to exploitation by LAAs. By compromising and intentionally manipulating the demand and supply of critical loads, attackers can rapidly alter the supply-demand balance, triggering dramatic price swings and seeking economic profits.
@@ III-B1 Mechanism of Energy Market
The energy market functions as a competitive platform where electricity generators offer their production to meet demand, while electric utilities and other market participants purchase electricity through real-time and day-ahead markets. In the day-ahead market, participants submit bids during the day-ahead bid period, followed by market clearing, where bids are processed, and energy schedules are set [87]. In the intra-day market, participants refine their bids and adjust their positions in response to updated information. The real-time market focuses on real-time flexibility bids, allowing providers to adjust supply and demand [88]. It balances the gaps between day-ahead commitments and the real-time electricity supply and demand in real time, which ensures grid stability by aligning supply with real-time demand.
Electricity pricing is a fundamental aspect of energy markets. Common pricing schemes include flat pricing, time-of-use (ToU) pricing, and real-time pricing (RTP). While flat pricing sets a constant price throughout the day, ToU pricing categorizes the day into different time periods, each with a specific price that is predetermined and announced before the operation begins [89]. In contrast, RTP adjusts electricity costs based on actual market conditions in real-time, typically on an hourly or sub-hourly basis, reflecting fluctuations in supply and demand. This approach factors in grid conditions, resource availability, and consumption changes. By exposing consumers to real-time price variations, RTP promotes flexible consumption behaviors, which enhances load management and grid stability. It is especially beneficial with the integration of RES, as it aligns consumption with periods of high generation or lower demand, ensuring efficient grid use [90].
@@ III-B2 LAAs to Disrupt the Energy Market Operation
Real-time electricity prices in energy markets are closely tied to system load demands [91], making them susceptible to manipulation through LAAs. By compromising large-scale loads, attackers can influence prices or disrupt the market [73, 74, 75, 58]. Understanding the impact of LAAs on the energy markets is crucial, as several studies demonstrate the feasibility of such attacks. Table IV summarizes existing research, which is further discussed below.
TABLE IV: Literature of LAAs effect in energy market operations.
@@
The manipulation of market via IoT (MaMIoT) attack, introduced in [75], leverages LAAs on high-wattage IoT botnets to affect the electricity prices in the energy market. An optimization model is proposed to solve for a stealthy and repeatable LAA, which maximizes the attack’s impact. By increasing the power demand of the compromised devices, the RTPs can be intentionally increased, which in turn enables power generation utilities to gain more benefits. Results demonstrate that even with a small number of controllable bots, the attacker can significantly increase their profits while inflicting substantial financial losses on other market participants.
The authors in [73] studied how LAAs affect locational marginal prices (LMPs) in distribution systems, showing that LAAs can propagate from targeted distribution systems through the transmission system to neighboring areas. Their findings reveal that LAAs at specific loads can substantially raise electricity prices locally. Moreover, energy storage system (ESS) owners can leverage LAAs to create high-loading conditions, enabling them to sell energy at inflated prices for profit. Similarly, Barreto et al. [74] examined demand response vulnerabilities, identifying two attacker types: malicious actors aiming to damage grid equipment and selfish attackers seeking profit. Results indicate that exploiting direct load control can significantly boost attacker profits.
The work in [58] investigated the charging manipulation attacks (CMAs) against EV charging, which aims to shift the EV aggregator’s demand to different periods of the day. The study delved into the impact of CMAs on aggregators’ profit and performance in the day-ahead markets and real-time markets by modeling their participation in both markets. Results show that the CMAs can lead to a considerable increase in the aggregated surcharge charging cost.
@@ IV Attack Detection and Localization
TABLE V: Detection and localization approaches against LAAs.
@@
If a successful LAA occurs, swift detection and localization are of utmost importance to limit the damage. Table V summarizes the work in this area, classifying approaches into model-based, data-driven, and hybrid methods. Model-based approaches utilize predefined algorithms or mathematical models to detect and locate LAAs, relying on a thorough understanding of system dynamics. Data-driven approaches utilize ML algorithms to automatically capture potential relationships in the data and anomalies caused by LAAs. Hybrid approaches use a combination of the two.
@@ IV-A Model-Based Approaches
One approach to detect LAAs is to model the power network as a semi-explicit class of differential-algebraic equations (DAEs) [92] and then use a sliding mode observer (SMO) to estimate the state of the system. The observer is able to estimate any signals representing failures to detect LAAs, with only the generator phase angles being measured. Focusing on DLAAs, [93] utilized the estimation of residual signals to detect attacks. Specifically, the power system was first modeled as state-space equations, and then a robust SMO was designed with an observer residual evaluation function. By comparing the value of this function to those observed in non-attack scenarios, the algorithm effectively identifies the DLAA.
Based on similar ideas of utilizing the observers, the sliding mode algorithm (STA) was used in the design of LAA detection while also incorporating the usage of ESS in the power system to defend against LAAs [94]. Another study utilized the linear matrix inequality (LMI) to design a bank of residual generators for functional observers to detect LAAs [95]. This approach reduces the number of observed states, making it suitable for large-scale power network calculations. The researchers in [50] designed a closed-loop SMO for the detection of closed-loop DLAAs, whose robustness is enhanced by using the $H_{\infty}$ performance index. A method using a low-rank Kalman filter was proposed in [96], which identifies destabilizing attacks by estimating system states and attack parameters with the rank-1 method, achieving high accuracy and computational efficiency.
Model-based approaches rely on observers or state estimators built using a state-space system model. While these methods provide strong theoretical foundations and require less data when the system is accurately modeled, the increasing integration of diverse equipment into power grids makes accurate modeling more challenging. This issue is particularly evident in large, complex grids, where developing precise state-space models requires significant effort, underscoring a key limitation of model-based approaches.